LadyRavyRay:

Use the formula V = Bh to enter a different version of this formula that you would use to find the area of the base B of a rectangular prism if you know the height h and the volume V. Complete the explanation for finding the equation. Begin with V = Bh and (select) both sides of the equation by (select) to get B =

LadyRavyRay:

@Astrid1 @AZ @nettym @jhonyy9 @snowflake0531

Florisalreadytaken:

you just have to solve for B ?

Florisalreadytaken:



Florisalreadytaken:

we need the 2nd one -- B=V / h

LadyRavyRay:

So the number answer is B = V/h

Florisalreadytaken:

that's just the formula

Florisalreadytaken:

fi the h and V is given, then yeah, we can find the B

Florisalreadytaken:

if h and V*

LadyRavyRay:

But, I'm not finding the answer.

AZ:

so we have V = Bh how do we move h to the other side to get B all by itself? h is being multiplied with B. So we need to do the opposite of that on both sides so we can move h to the other side. What is the opposite of multiplication?

LadyRavyRay:

Division

AZ:

Good and what are we dividing both sides by? (Hint: re-read what I wrote, I already told you.. but basically we're dividing by the thing that's there with B so we can get B all by itself on one side)

LadyRavyRay:

H or V?

AZ:

V = B*h what is B being multiplied with? We need to divide that on both sides

LadyRavyRay:

h

AZ:

Exactly! And what do you get when divide both sides by 'h'

LadyRavyRay:

B = V/h?

AZ:

Wonderful! And now you should be able to fill in the sentence We've already said what goes in the blank you're dividing by h on both sides to get that equation

LadyRavyRay:

Begin with V = Bh and divide both sides of the equation by h to get B = V/h.

AZ:

There you have it!

LadyRavyRay:

Thank you. I have another question too.

AZ:

Sure, tag me in your next post :)

LadyRavyRay:

Okie.